### v1

# RWKV-CLIP: A Robust Vision-Language Representation Learner

Abstract
Contrastive Language-Image Pre-training (CLIP) has significantly improved performance in various vision-language tasks by expanding the dataset with image-text pairs obtained from websites. This paper further explores CLIP from the perspectives of data and model architecture. To address the prevalence of noisy data and enhance the quality of large-scale image-text data crawled from the internet, we introduce a diverse description generation framework that can leverage Large Language Models (LLMs) to synthesize and refine content from web-based texts, synthetic captions, and detection tags. Furthermore, we propose RWKV-CLIP, the first RWKV-driven vision-language representation learning model that combines the effective parallel training of transformers with the efficient inference of RNNs. Comprehensive experiments across various model scales and pre-training datasets demonstrate that RWKV-CLIP is a robust and efficient vision-language representation learner, it achieves state-of-the-art performance in several downstream tasks, including linear probe, zero-shot classification, and zero-shot image-text retrieval. To facilitate future research, the code and pre-trained models are released at https://github.com/deepglint/RWKV-CLIP * Equal contribution. $\dagger$ Corresponding author.
1 Introduction
The proliferation of mobile networks and social platforms has greatly accelerated the large-scale production of image-text pairs. This unprecedented abundance of data has established the foundation for vision-language pre-training. Contrastive Language-Image Pre-training (CLIP) employs two distinct unimodal encoders for images and text, utilizing a contrastive loss, a highly effective mechanism for representation learning. Having been pre-trained on extensive image-text pairs collected from the internet, CLIP demonstrates strong transferability and has been widely applied across various domains Zhou et al. (2023); Yao et al. (2023).
Figure 1: The proposed RWKV-CLIP combines the effective parallel training of transformers with the efficient inference of RNNs, achieving better efficiency and accuracy than the baseline methods (e.g., CLIP and ALIP).
In recent years, many large-scale image-text datasets collected from the internet have been released. LAION400M Schuhmann et al. (2021) is created for research purposes and it contains 400 million image-text pairs curated using the CLIP model. Building on this, LAION5B Schuhmann et al. (2022), which consists of 5.85 billion CLIP-filtered image-text pairs, successfully replicates and fine-tunes basic models such as CLIP. However, using the CLIP model to filter web-based image-text pairs still retains a considerable presence of noisy data. To improve data quality, DataComp Gadre et al. (2024) employs various strategies such as basic filtering, CLIP score filtering, and text&image-based filtering. However, inherent characteristics of internet data, such as abstract text representations and semantic discrepancies between text and images, remain significant obstacles.
In recent years, the Transformer Vaswani et al. (2017) model has been extensively applied in large-scale representation learning, yielding significant performance improvements across multiple downstream tasks Acosta et al. (2022); Kirillov et al. (2023); Wang et al. (2023b), including image classification Dosovitskiy et al. (2020); Wang et al. (2023a), text generation Brown et al. (2020), and speech recognition Radford et al. (2023). Despite these achievements, the quadratic computational complexity inherent in Transformer limits its capacity to effectively process high-resolution images and long sequences, posing a substantial challenge to its broader applicability across varied domains.
In this paper, we design a framework for generating diverse descriptions. Following ALIP Yang et al. (2023), we first use the OFA Wang et al. (2022) model to generate synthetic descriptions consistent with image content. However, constrained by the training data, OFA can only partially identify coarse-grained object categories. Therefore, we introduce an open-set image tagging model RAM++ Huang et al. (2023) to capture more detailed and precise semantic information from images. By leveraging LLMs, we synthesize and refine information from web-based texts, synthetic captions, and detection tags. Additionally, inspired by RWKV Peng et al. (2024) and Vision-RWKV Duan et al. (2024), we propose RWKV-CLIP, the first RWKV-driven vision-language representation learning model that combines the effective parallel training of Transformers with the efficient inference of RNNs. Extensive experiments across various model scales and pre-training datasets demonstrate that RWKV-CLIP is a robust and efficient vision-language representation learner. The main contributions of this paper are summarized as follows:
  • We introduce a diverse description generation framework, which can leverage LLMs to synthesize and refine information from web-based texts, synthetic captions, and detection tags to produce more accurate and semantically enriched descriptions.
  • We propose the RWKV-CLIP, the first RWKV-driven vision-language representation learning model, which combines the parallel training effectiveness of Transformers with the inference efficiency of RNNs.
  • We demonstrate the robustness and effectiveness of RWKV-CLIP as a vision-language representation learner through extensive experiments across various model scales and pre-training datasets.
2 Related Work
Figure 2: The architecture of our proposed diverse description generation framework.
2.1 Vision-Language Representation Learning
As the milestone in vision-language representation learning, CLIP Radford et al. (2021) has garnered unparalleled interest due to its remarkable zero-shot recognition capability and outstanding transfer performance. Subsequently, a significant amount of enhancement works based on CLIP have been proposed. SLIP Mu et al. (2022) combines self-supervised learning with CLIP pre-training to achieve significant performance improvements. DeCLIP Li et al. (2022b) employs multi-view supervision across modalities and nearest-neighbor supervision from similar pairs to enhance representation learning efficiency. FILIP Yao et al. (2022) refines contrastive loss to learn fine-grained representations for image patches and sentence words. UniCLIP Lee et al. (2022) boosts data efficiency by integrating contrastive loss across multiple domains into a single universal space. HiCLIP Geng et al. (2023) enhances cross-modal alignment by incorporating hierarchy-aware attention into both visual and language branches of CLIP. ALIP Yang et al. (2023) introduces a gating mechanism to reduce the influence of noisy pairs using synthetic data. Different from the above methods, this paper conducts further exploration of both the data and model architecture, proposing a diverse description generation framework and introducing RWKV-CLIP, the first RWKV-driven vision-language representation model.
2.2 Text Augumentation
With the success of LLMs in Natural Language Processing (NLP), there is growing interest in leveraging LLMs to enhance text descriptions in large-scale image-text pairs. LaCLIP Fan et al. (2023) explores different strategies to generate rewrite examples and uses the in-context learning ability of LLMs to rewrite text within image-text datasets. However, the hallucination issue of LLMs and reliance on limited samples to guide the rewriting process can still introduce significant noise. To address this, CapsFusion Yu et al. (2024) generates synthetic captions for each image and utilizes ChatGPT to merge raw texts and synthetic captions, creating a dataset with one million instructions for LLaMA fine-tuning. Despite this, caption generation models such as OFA Wang et al. (2022) and BLIP Li et al. (2022a) are limited by their training data and can only identify a restricted set of coarse-grained object categories. In this paper, we introduce the open-set image tagging model RAM++ Huang et al. (2023) to assign semantic detection tags to each image. Beneficial from detection tags, more semantic information can be introduced from images, which in turn further constrains LLMs and mitigates hallucinations.
2.3 Receptance Weighted Key Value
RWKV Peng et al. (2023) is first proposed in NLP, it addresses memory bottleneck and quadratic scaling in Transformers through efficient linear scaling while retaining expressive characteristics like parallelized training and robust scalability. Recently, Vision-RWKV Duan et al. (2024) successfully transferred the RWKV from NLP to vision tasks, outperforming ViT in image classification with faster processing and reduced memory consumption for high-resolution inputs. PointRWKV He et al. (2024) demonstrates leading performance across various downstream tasks, surpassing Transformer- and Mamba-based counterparts in efficiency and computational complexity. Furthermore, Diffusion-RWKV Fei et al. (2024) adapts RWKV for diffusion models in image generation tasks, achieving competitive or superior performance compared to existing CNN or Transformer-based diffusion models. However, these methods have only validated RWKV in specific downstream tasks, and the potential of RWKVs to replace ViTs in vision-language representation learning remains unverified.
3 Method
Figure 3: The architecture of RWKV-CLIP, which consists of M×$\times$ and N×$\times$ RKWV-driven blocks followed by an average pooling layer.
In this section, we first introduce a diverse description generation framework that leverages the capabilities of large language models to integrate information from web-based texts, synthetic captions, and detection tags. Subsequently, we provide a detailed exposition of RWKV-CLIP.
3.1 Diverse Description Generation
The architecture of our proposed diverse description generation framework is illustrated in Fig. 2. To mitigate the effects of mismatched image-text pairs, following ALIP Yang et al. (2023), we first adopt the OFAbase model to generate a synthetic caption for each image. The synthetic captions exhibit a high degree of semantic alignment with the image, facilitating alignment across different modal feature spaces. However, constrained by the training data, OFAbase can recognize a limited number of object categories and tends to produce captions with a simplistic sentence structure. To capture finer-grained semantic information within images, we incorporate the open-set image tagging models RAM++ Huang et al. (2023) to extract object detection tags for each image.
To assess the viability of our approach, following CapsFusion Yu et al. (2024), we initially leverage ChatGPT to combine information from raw texts, synthetic captions, and detection tags. However, the time and computational effort involved is prohibitive. Therefore, we construct an instruction dataset based on ChatGPT interactions and fine-tuned the open-source LLaMA3 with this dataset. After that, we leverage the fine-tuned LLaMA3 model for large-scale inference. Specifically, we select 70K image-text pairs from YFCC15M that have more than 10 detection tags. Then, we input the raw texts, synthetic captions, and detection tags of these data into ChatGPT to get instruction responses. The details of the instruction prompt are provided in the supplementary material.
After obtaining the instruction dataset, we utilize the LLaMA Factory Zheng et al. (2024) to finetune the LLaMA3-8B and leverage vLLM Kwon et al. (2023) to accelerate large-scale inference.
3.2 RWKV-CLIP
In this section, we propose RWKV-CLIP, a robust and efficient RWKV-driven vision-language representation learner. Inspired by CLIP Radford et al. (2021) and Vision-RWKV Duan et al. (2024), RWKV-CLIP adopts a dual-tower architecture with a block-stacked encoder design like the Transformer Vaswani et al. (2017), where each block consists of a spatial mixing and a channel mixing module. The overview architecture of our proposed RWKV-CLIP is shown in Fig. 3.
Input Augmentation. Based on our proposed diverse description generation framework, we can obtain three types of text: raw text $T_{r}$, synthetic caption $T_{s}$, and generated description $T_{g}$. To improve the robustness of the model, we randomly select a text from $[T_{r},T_{s},T_{g}]$ as the augmentation for text inputs:
$\mathrm{aug}(T)=\mathrm{Sample}([T_{r},T_{s},T_{g}])$ (1)
Meanwhile, the input image $I\in\mathbb{R}^{H\times W\times 3}$ is transformed into $HW/p^{2}$ patches, where $p$ is the patch size.
Spatial Mixing. The input text $\text{aug}(T)$ and image $I$ are passed through the spatial mixing module, which acts as an attention mechanism and performs global attention computation of linear complexity. Specifically, the input data is shifted and entered into four parallel linear layers to obtain multi-head vectors $G_{x}^{s},R_{x}^{s},K_{x}^{s},V_{x}^{s}$:
$\psi_{x}^{s}=\mathrm{Lerp}_{\psi}(x)\cdot w_{\psi}^{s},\quad\psi\in\{G,R,K,V\},$ (2)
where $\rm Lerp$ is the linear interpolation Peng et al. (2024). In this paper, we adopt Q-Lerp and B-Lerp for image and text encoders respectively. The Q-Lerp can be formulated as:
[…]
$I^{\star}=\text{Concat}(I_{1},I_{2},I_{3},I_{4}).$
The B-Lerp can be presented as:
[…]
$T^{\star}=\text{Concat}(T_{1},T_{2}),$
where $\Psi\in\{G,R,K,V,w\}$, $\eta_{\Psi}$ denotes learnable vectors, $I^{\star}$ is the quad-directional shift vector in the image, i.e., $I_{1}=x[h-1,w,0:C/4],I_{2}=x[h+1,w,C/4:C/2],I_{3}=x[h,w-1,C/2:3C/4],I_{4}=x[h,w+1,3C/4:C]$, $T^{\star}$ is the bi-directional shift in the text i.e., $T_{1}=[w-1,0:C/2]$, $T_{2}=[w+1,C/2:C]$, where $h,w,C$ present the number of height, width, and channel. These shift functions enhance feature interaction at the channel level, enabling a focus on neighboring tokens. Specifically, the bi-directional shift ensures forward and backward interaction of text tokens without increasing additional FLOPs. To avoid a fixed learned vector, a new time-varying decay $w_{x}$ is calculated as follows:
$\phi(x)=\lambda+\tanh(x\cdot M_{i})\cdot M_{j},$ (5)
$\hat{w}_{x}^{s}=x+(1-\phi(\text{Lerp}_{w}(x)))\cdot x^{\star},$
$\tilde{w}_{x}^{s}=\phi(\hat{w}_{x}^{s}),w_{x}^{s}=\exp\left(-\exp(\tilde{w}_{x}^{s})\right),$
where $x\in\{I,T\}$, $\lambda$ is a learnable vector, $M_{i},M_{j}$ are learnable weight matrices. The function $\phi$ is used to obtain learned vectors by inexpensively augmenting inputs with additional offsets. $\hat{w}_{x}^{s}$ and $\tilde{w}_{x}^{s}$ are middle values of $w_{x}^{s}$ during the calculation process. This process allows each channel of $w_{x}$ to vary based on a mix of the current and prior tokens $x^{\star}$.
Subsequently, $w_{x}^{s},R_{x}^{s},K_{x}^{s},V_{x}^{s}$ are used to compute the global attention result $wkv_{t}$ via a linear complexity bidirectional attention mechanism. This result is then multiplied by $\sigma(G_{x}^{s})$, functioning as a gate mechanism to control the output $O_{x}^{s}$:
$wkv_{t}=\text{Bi-WKV}_{t}(w_{x}^{s},R_{x}^{s},K_{x}^{s},V_{x}^{s}),$ (6)
$O_{x}^{s}=\text{Concat}\left(\sigma(G_{x}^{s})\odot\text{LN}(wkv_{t})\right)\cdot w_{o}^{s},$
where $\sigma(\cdot)$ denotes the SiLU function Elfwing et al. (2018), and $\odot$ means element-wise multiplication, $\rm LN$ is the layer norm and the Bi-WKV Duan et al. (2024); Peng et al. (2024) can be formulated as:
[…]
$\quad+\sum_{i=0}^{t-1}\text{diag}(\epsilon_{i,j})\cdot K_{s,i}^{\mathsf{T}}\cdot V_{s,i}$
$\quad+\sum_{i=t+1}^{T-1}\text{diag}(\epsilon_{i,j})\cdot K_{s,i}^{\mathsf{T}}\cdot V_{s,i})$
where $u$ is a per-channel learned boost and $\epsilon_{i,j}=\odot^{i-1}_{j=1}w_{j}$ is a dynamic decay.
Channel Mixing. The spatial mixing module is followed by the channel-mixing module. Similarly, the $R_{x}^{c},K_{x}^{c}$ are obtained by Lerp:
$\psi_{x}^{c}=\text{Lerp}_{\psi}(x)\cdot w_{\psi}^{c},\quad\psi\in\{R,K\}$ (8)
After that, a linear projection and a gate mechanism are performed respectively and the final output $O_{x}^{c}$ is formulated as:
$O_{x}^{c}=(\sigma(R_{x}^{c})\odot\rho(K_{x}^{c}))\cdot w_{o}^{c}$ (9)
where $\rho$ is the squaredReLU Agarap (2018). After passing through the stack RWKV-based image and text encoders $E_{I}$ and $E_{T}$, we can get the image embeddings $\hat{I}=E_{I}(I)$ and text embeddings $\hat{T}=E_{T}(aug(T))$, the loss function $L$ is defined as:
$\begin{split}L=-\sum_{i=1}^{N}\left[\log\frac{e^{\hat{I}_{i}^{\top}\hat{T}_{i}/\tau}}{\sum_{j}e^{\hat{I}_{i}^{\top}\hat{T}_{j}/\tau}}+\log\frac{e^{\hat{I}_{i}^{\top}\hat{T}_{i}/\tau}}{\sum_{j}e^{\hat{I}_{j}^{\top}\hat{T}_{i}/\tau}}\right]\end{split}$ (10)
4 Experiments
Method Pre-train dataset
CIFAR10
CIFAR100
Food101
Pets
Flowers
SUN397
Cars
DTD
Caltech101
Aircraft
Average
CLIP-ViT-B/32Radford et al. (2021) YFCC15M 86.5 64.7 69.2 64.6 90.6 66.0 24.9 61.3 79.1 23.1 63.0
DeCLIP-ViT-B/32 Li et al. (2022b) YFCC15M 89.2 69.0 75.4 72.2 94.4 71.6 31.0 68.8 87.9 27.6 68.7
HiCLIP-ViT-B/32 Geng et al. (2023) YFCC15M 89.5 71.1 73.5 70.6 91.9 68.8 30.8 63.9 84.8 27.4 67.2
ALIP-ViT-B/32 Yang et al. (2023) YFCC15M 94.3 77.8 75.8 76.0 95.1 73.3 33.6 71.7 88.5 36.1 72.2
RWKV-CLIP-B/32 YFCC15M 95.3 81.8 76.4 77.1 92.4 73.1 37.7 73.2 90.6 43.5 74.1
Table 1: Linear probe performance on 10 downstream datasets. RWKV-CLIP achieves an average performance improvement of 1.9$\sim$11.1%.
Text retrieval Image retrieval
Flickr30k MSCOCO Flickr30k MSCOCO
Method R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP-ViT-B/32Radford et al. (2021) 34.9 63.9 75.9 20.8 43.9 55.7 23.4 47.2 58.9 13.0 31.7 42.7
SLIP-ViT-B/32 Mu et al. (2022) 47.8 76.5 85.9 27.7 52.6 63.9 32.3 58.7 68.8 18.2 39.2 51.0
DeCLIP-ViT-B/32 Li et al. (2022b) 51.4 80.2 88.9 28.3 53.2 64.5 34.3 60.3 70.7 18.4 39.6 51.4
UniCLIP-ViT-B/32 Lee et al. (2022) 52.3 81.6 89.0 32.0 57.7 69.2 34.8 62.0 72.0 20.2 43.2 54.4
HiCLIP-ViT-B/32 Geng et al. (2023) - - - 34.2 60.3 70.9 - - - 20.6 43.8 55.3
ALIP-ViT-B/32 Yang et al. (2023) 70.5 91.9 95.7 46.8 72.4 81.8 48.9 75.1 82.9 29.3 54.4 65.4
RWKV-CLIP-B/32 76.0 94.7 97.6 50.3 76.2 85.2 57.6 82.3 88.7 34.0 60.9 71.7
Table 2: Zero-shot image-text retrieval performance on the test splits of Flickr30k and MSCOCO. RWKV-CLIP achieves a significant improvement on all metrics.
4.1 Experimental Settings
Pre-training Datasets. We train our model on the YFCC15M dataset, which is a subset of YFCC100M Thomee et al. (2016) filtered by DeCLIP Li et al. (2022b). To further verify the effectiveness and generalizability of RWKV-CLIP, following ALIP Yang et al. (2023), we randomly select subsets of 10M and 30M from the LAION400M Schuhmann et al. (2021). We then conduct a series of experiments with different model scales and pre-training datasets.
Implementation Details. Consistent with ALIP Yang et al. (2023), we employ OFAbase to generate synthetic captions. The instruction dataset is constructed using ChatGPT-35-turbo, and we fine-tune LLaMA3-8B to enhance the generation of diverse descriptions. We employ AdamW Loshchilov and Hutter (2019) as the optimizer, initialized with a learning rate of $31e-$ 3 and a weight decay of $0.2$. The parameters $\beta{1}$ and $\beta_{2}$ are set to $0.9$ and $0.98$, respectively. The input image size is $224\times 224$, and the input text sequence length is truncated or padded to $77$. The temperature parameter $\tau$ is initialized to $0.07$. We train RWKV-CLIP for 32 epochs with a batch size of $4096$ on $8$ NVIDIA A100(80G) GPUs. We meticulously regulate the parameters and FLOPs of RWKV-CLIP to ensure the fairness of the experimental comparison. Please refer to the supplementary material for more detailed parameters, FLOPs, and settings of RWKV-CLIP.
4.2 Experimental Results
Linear Probe. Building upon previous works Yang et al. (2023); Li et al. (2022b); Geng et al. (2023), we use RWKV-CLIP as a feature extractor and train only a logistic regression classifier. Tab. 1 details the linear probe performance across 10 downstream datasets, as referenced in ALIP Yang et al. (2023). RWKV-CLIP achieves a significant performance improvement ranging from 1.9% $\sim$ 11.1% over the baseline models, outperforming ALIP in 8 of the 10 datasets. The observed performance improvements are primarily due to two main factors: (1) Our proposed description generation framework effectively synthesizes and refines information from web-based texts, synthetic captions, and detection tags, producing more accurate and semanti- cally enriched descriptions. (2) RWKV-CLIP exhibits superior representation learning capabilities compared to ViT-based models.
Zero-shot Image-text Retrieval. In Tab. 2, we compare our method with state-of-the-art approaches in zero-shot image-text retrieval on Flickr30k and MSCOCO. RWKV-CLIP achieves new state-of-the-art results on all evaluation metrics. Specifically, RWKV-CLIP achieves 76.0%/57.6% I2T/T2I retrieval Recall@1 on Flickr30K, surpassing ALIP by 5.5%/8.7%. Similarly, significant improvements of 3.5%/4.7% in I2T/T2I retrieval Recall@1 are observed for RWKV-CLIP on MSCOCO. This exceptional image-text retrieval capability indicates that the representations learned by RWKV-CLIP are robust and exhibit enhanced cross-modal alignment.
Method Pre-train dataset
CIFAR10
CIFAR100
Food101
Pets
Flowers
SUN397
Cars
DTD
Caltech101
Aircraft
ImageNet
Average
CLIP-ViT-B/32Radford et al. (2021) YFCC15M 63.7 33.2 34.6 20.1 50.1 35.7 2.6 15.5 59.9 1.2 32.8 31.8
SLIP-ViT-B/32 Mu et al. (2022) YFCC15M 50.7 25.5 33.3 23.5 49.0 34.7 2.8 14.4 59.9 1.7 34.3 30.0
FILIP-ViT-B/32 Yao et al. (2022) YFCC15M 65.5 33.5 43.1 24.1 52.7 50.7 3.3 24.3 68.8 3.2 39.5 37.2
DeCLIP-ViT-B/32 Li et al. (2022b) YFCC15M 66.7 38.7 52.5 33.8 60.8 50.3 3.8 27.7 74.7 2.1 43.2 41.3
HiCLIP-ViT-B/32 Geng et al. (2023) YFCC15M 74.1 46.0 51.2 37.8 60.9 50.6 4.5 23.1 67.4 3.6 40.5 41.8
ALIP-ViT-B/32 Yang et al. (2023) YFCC15M 83.8 51.9 45.4 30.7 54.8 47.8 3.4 23.2 74.1 2.7 40.3 41.7
RWKV-CLIP-B/32 YFCC15M 79.8 55.1 50.6 37.6 57.1 54.0 4.1 24.6 77.1 4.0 44.3 44.4
Table 3: Zero-shot classification performance on 11 downstream datasets. RWKV-CLIP achieves an average performance improvement of 2.6$\sim$12.6%.
Zero-shot Classification. We present the zero-shot classification performance across 11 datasets. To ensure fair comparisons, we use the same prompt templates and class names as established in ALIP Yang et al. (2023) and SLIP Mu et al. (2022). As shown in Tab. 3, RWKV-CLIP achieves an average performance improvement of 2.6% $\sim$ 14.4% over baseline models. Notably, our model outperforms ALIP in 10 out of the 11 datasets, with significant enhancements on instance discrimination datasets such as Food101, and ImageNet. This improvement is mainly due to the diverse descriptions generated by our framework, providing more fine-grained semantic information.
Zero-Shot Robustness Evaluation. In Tab. 4, we present a robustness evaluation comparing ALIP and RWKV-CLIP. Our results show that RWKV-CLIP consistently outperforms ALIP in terms of robustness across all datasets with an average improvement of 2.0%. These experimental results establish the RWKV-driven model as a robust representation learner.
Method IN-V2 IN-A IN-R IN-Sketch Average
ALIP-ViT-B/32 34.1 16.1 35.2 12.1 24.4
RWKV-CLIP-B/32 37.5 16.7 37.0 14.5 26.4
Table 4: Zero-shot robustness comparison of ALIP and RWKV-CLIP pretrained on YFCC15M.
4.3 Ablation Study
Effectiveness of Model and Data Scaling. To evaluate the effectiveness of RWKV-CLIP on model and data scaling, we conduct experiments on randomly selected subsets of 10M and 30M from LAION400M. For a more comprehensive comparison, we report the linear probe performance on 26 downstream datasets. As shown in Fig. 5, RWKV-CLIP significantly improves performance across different model scales and pre-training datasets. These results demonstrate the robustness and extensibility of RWKV-CLIP. Detailed experimental results can be found in the supplementary material.
Comparision Analysis with CapsFusion. To further demonstrate the performance differences between our proposed diverse description generation framework and CapsFusion, we used CapsFusion-LLaMA to rewrite the YFCC15M dataset based on raw texts and synthetic captions. We then trained RWKV-CLIP using texts generated by our framework and CapsFusion. As shown in Tab. 5, our framework achieves a 0.9% and 2.1% improvement in the average linear probe and zero-shot classification performance, respectively. This improvement is primarily due to the detection tags introducing more semantic information from images, which further constrains LLMs and reduces hallucinations (as shown in Fig. 4).
Method Text Generation Model Linear probe Avg Zero-shot Avg
RWKV-CLIP-B/32 CapsFusion 72.6 33.1
RWKV-CLIP-B/32 Ours 73.5 35.2
Table 5: Performance comparison using text generated by our proposed diverse description generation framework vs. CapsFusion.
Figure 4: Comparison of our proposed diverse description generation framework vs. CapsFusion. Hallucinations are highlighted in red, and additional semantic information is highlighted in green.
Figure 5: Linear probe performance comparison between RWKV-CLIP and ALIP on 26 downstream datasets. The comparisons include RWKV-CLIP-B/32 vs. ALIP-ViT-B/32 on LAION10M, RWKV-CLIP-B/16 vs. ALIP-ViT-B/16 on LAION10M, and RWKV-CLIP-B/32 vs. ALIP-ViT-B/32 on LAION30M, presented from left to right.
$T_{r}$ $T_{s}$ $T_{g}$ Dataset Linear probe Avg Zero-shot Avg
YFCC15M 71.3 38.7
YFCC15M 72.4 23.1
YFCC15M 73.5 35.2
YFCC15M 73.0 43.0
YFCC15M 73.8 43.4
YFCC15M 74.1 44.4
Table 6: Ablation experiment results using different types of text. $T_{r}$: raw text. $T_{s}$: synthetic caption. $T_{g}$: generated diverse description using our proposed framework.
Ablation on Different Types of Text. We conduct ablation experiments on different categories of text, the average linear probe results on 10 datasets and the average zero-shot classification accuracy on 11 datasets are shown in Tab. 6. Synthetic captions and generated diverse descriptions yielded superior linear probe performance compared to raw texts. This improvement is attributed to the high incidence of mismatched image-text pairs in raw texts, which can adversely affect representation learning. As shown in Fig. 6, our analysis of cosine similarity (computed by CLIP-L14) and token counts across different text types reveals that synthetic captions and generated diverse descriptions have higher average similarity and token counts than raw texts. Furthermore, despite these advantages, raw texts achieve superior zero-shot classification results, mainly due to the constraints imposed by the prompt template.
Figure 6: Statistical analysis of raw texts, synthetic captions, and generated diverse descriptions on the YFCC15M.
4.4 Ablation on Model Architecture
In Tab. 7, base on text augmentation, we perform an ablation study combining RWKV and Transformer architectures. Compared with TransformerI and TransformerT, the integration of RWKVI and TransformerT achieves a 2.7% improvement on linear probe but the zero-shot classification performance declines by 10.8%. This reduction is primarily due to the poor compatibility between the RWKV and Transformer architectures. Conversely, the combination of RWKVI and RWKVT yields improvements of 3.2% and 2.7% in linear probe and zero-shot classification, respectively, indicating that RWKV outperforms Transformer in vision-language representation learning.
Image Text Linear Probe Zero-shot
RWKVI TransformerI RWKVT TransformerT Avg Avg
70.9 41.7
73.6 30.9
71.0 41.1
74.1 44.4
Table 7: Ablation on model architecture.
Analysis of Feature Embedding. To understand what makes RWKV-CLIP effective, we randomly select 250 image-text pairs from YFCC15M and visualize the modality gaps of ALIP and RWKV-CLIP. Specifically, each image and its corresponding text are encoded into embedding space and reduced to two dimensions using UMAP McInnes et al. (2018). As shown in Fig. 7, we found that the representations learned by RWKV-CLIP exhibit clearer discriminability within the same modality. Additionally, compared to ALIP, RWKV-CLIP demonstrates closer distances in the image-text modality space, indicating superior cross-modal alignment performance.
Figure 7: Visualization of modality gaps.
5 Conclusion
In this paper, we further explore CLIP from the perspectives of data and model architecture. We introduce a diverse description generation framework that can leverage Large Language Models (LLMs) to combine and refine information from web-based image-text pairs, synthetic captions, and detection tags. Besides, we propose RWKV-CLIP, the first RWKV-driven vision-language representation learning model that combines the effective parallel training of transformers with the efficient inference of RNNs. Our method demonstrates superior performance across various model scales and pre-training datasets on different downstream tasks. We hope that our work provides insights into vision-language representation learning models.
Appendix A Detail Experimental Settings
A.1 Model Architectures
We meticulously regulate the parameters and FLOPs of RWKV-CLIP to ensure the fairness of the experimental comparison. The detailed parameters and FLOPs of RWKV-CLIP-B/32 and RWKV-CLIP-B/16 are shown in Tab. 8. The detail settings of RWKV-CLIP-B/32 and RWKV-CLIP-B/16 are shown in Tab. 10.
Method Image Text Total
Params(M) FLOPs(G) Params(M) FLOPs(G) Params(M) FLOPs(G)
CLIP-ViT-B/32 87.85 8.73 63.44 5.82 151.29 14.55
RWKV-CLIP-B/32 84.21 7.91 65.35 4.93 149.56 12.84
CLIP-ViT-B/16 86.19 33.72 63.44 5.82 149.63 39.54
RWKV-CLIP-B/16 82.83 31.05 65.35 4.93 148.18 35.98
Table 8: Parameters and FLOPs comparison between CLIP and RWKV-CLIP.
A.2 Detail Instruction Prompt
The prompt used to input ChatGPT is present in the following:
"Please merge the information from the given raw text and the synthetic caption with the help of the highly relevant detection tags. The raw caption offers detailed real-world information, yet it suffers from flaws in sentence structure and grammar. The synthetic caption exhibits impeccable sentence structure but often lacks in-depth real-world details and may contain false information. The highly relevant detection tags are provided to enrich the semantic information of the raw caption, while some are redundant and noisy. You are a great information integration and summary expert, you are also good at enriching semantic information. Ensure a well-structured sentence while retaining the detailed real-world information provided in the raw caption. Avoid simply concatenating the sentences and avoid adding external information to describe. Correctness and simplify sentences finally. Raw caption:<raw caption>, synthetic caption:<synthetic caption>, and highly relevant detection tags:<detection tags>".
A.3 Experimental Settings
We present the settings used in the training RWKV-CLIP in Tab. 9.
Hyperparameter Value
Initial temperature $0.07$
Adam $\beta_{1}$ $0.9$
Adam $\beta_{2}$ $0.98$
Adam $\epsilon$ $10^{-6}$
Weight decay $0.2$
Batch size 4096
Learning rate 0.001
Learning rate scheduler OneCycleLR
Pct start 0.1
Training epochs 32
GPU 8×8$\times$8 ×A100
Table 9: Hyperparameters used for RWKV-CLIP pre-training.
A.4 Prompts for Zero-shot Classification
In this work, we evaluate the zero-shot performance of RWKV-CLIP on 11 downstream datasets. All the prompts for the 11 downstream datasets are presented in Tab. 13.
Embedding Input Image Encoder Text Encoder
Model dimension resolution layers hidden rate heads Init layers hidden rate heads Init
RWKV-CLIP-B/32 640 224 12 5 8 6 3.5 10
RWKV-CLIP-B/16 640 224 12 5 8 6 3.5 10
Table 10: The detail architecture parameters for our proposed RWKV-CLIP.
Method Pre-train data
Food101
CIFAR10
CIFAR100
Birdsnap
SUN397
Cars
Aircraft
VOC2007
DTD
Pets
Caltech101
Flowers
MNIST
STL10
EuroSAT
RESISC45
GTSRB
KITTI
Country211
PCAM
UCF101
Kinetics700
CLEVR
Memes
SST2
ImageNet
Average
ALIP-ViT-B/32 LAION10M 71.5 92.2 76.1 36.3 67.3 70.1 41.8 85.3 71.3 74.3 86.9 90.7 98.0 94.6 95.4 84.3 84.1 70.0 12.9 83.4 75.9 46.4 51.0 54.8 56.5 59.6 70.4
RWKV-CLIP-B/32 LAION10M 72.7 94.7 81.4 42.3 68.3 70.3 47.9 86.5 73.6 76.6 90.0 89.4 99.0 94.6 97.0 85.6 87.0 74.9 13.8 85.1 80.8 49.3 60.6 55.4 58.3 63.7 73.0
ALIP-ViT-B/16 LAION10M 77.2 93.3 77.0 45.1 69.4 77.3 48.6 87.7 74.5 79.0 88.1 93.0 98.3 96.3 96.3 86.4 83.7 72.2 14.2 85.2 80.1 50.1 55.4 55.7 57.3 64.8 73.3
RWKV-CLIP-B/16 LAION10M 78.9 95.1 81.8 50.3 72.0 76.8 50.3 89.4 75.4 79.7 91.9 91.7 99.0 96.4 96.9 87.8 87.4 75.7 15.2 85.5 83.9 53.0 61.8 55.9 60.0 68.4 75.4
ALIP-ViT-B/32 LAION30M 76.6 94.0 79.3 44.2 70.6 77.7 48.4 87.6 74.4 80.4 90.0 93.8 98.3 96.3 96.0 86.7 84.7 72.3 15.0 85.0 81.0 50.6 55.6 56.1 59.8 65.0 73.8
RWKV-CLIP-B/32 LAION30M 76.6 95.6 82.8 46.0 71.0 77.9 50.0 88.2 74.5 78.9 91.6 92.1 99.0 96.5 97.1 86.9 87.6 78.9 15.2 85.6 83.4 51.8 61.6 58.9 58.9 67.2 75.2
Table 11: Top-1 accuracy(%) of linear probe on 26 image classification datasets.
Appendix B Detail Linear Probe on LAION
Dataset Classes Train size Test size Evaluation metric
Food101 102 75,750 25,250 accuracy
CIFAR10 10 50,000 10,000 accuracy
CIFAR100 100 50,000 10,000 accuracy
Birdsnap 500 42,138 2,149 accuracy
SUN397 397 19,850 19,850 accuracy
Cars 196 8,144 8,041 accuracy
Aircraft 100 6,667 3,333 mean per class
VOC2007 20 5011 4952 11-point mAP
DTD 47 3,760 1,880 accuracy
Pets 37 3,680 3,669 mean per class
Caltech101 101 3,000 5,677 mean-per-class
Flowers 102 2,040 6,149 mean per class
MNIST 10 60,000 10,000 accuracy
STL10 10 5,000 8,000 accuracy
EuroSAT 10 10,000 5,000 accuracy
RESISC45 45 3,150 25,200 accuracy
GTSRB 43 26,640 12,630 accuracy
KITTI 4 6770 711 accuracy
Country211 211 42,200 21,100 accuracy
PCAM 2 294,912 32,768 accuracy
UCF101 101 9,537 1,794 accuracy
Kinetics700 700 530,779 33,944 mean(top1,top5)
CLEVR 8 2,000 500 accuracy
Memes 2 8,500 500 ROC AUC
SST2 2 7,792 1,821 accuracy
ImageNet 1000 1,281,167 50,000 accuracy
Table 12: List of linear probe datasets with the data distribution and evaluation metrics.
B.1 Downstream Datasets
To comprehensively demonstrate the performance of RWKV-CLIP, we compared the linear probe results of RWKV-CLIP and ALIP across 26 datasets. These datasets include Food101 Bossard et al. (2014), CIFAR10 Krizhevsky et al. (2009), CIFAR100 Krizhevsky et al. (2009), Birdsnap Berg et al. (2014), SUN397 Xiao et al. (2010), Stanford Cars Krause et al. (2013), FGVC Aircraft Maji et al. (2013), VOC2007 Everingham (2007), DTD Cimpoi et al. (2014), Pets Parkhi et al. (2012), Caltech101 Fei-Fei et al. (2004), Flowers102 Nilsback and Zisserman (2008), MNIST LeCun et al. (1998), SLT10 Coates et al. (2011), EuroSAT Helber et al. (2019), RESISC45 Cheng et al. (2017), GTSRB Stallkamp et al. (2012), KITTI Geiger et al. (2012), Country211 Radford et al. (2021), PCAM Veeling et al. (2018), UCF101 Soomro et al. (2012), Kinetics700 Carreira et al. (2019), CLEVR Johnson et al. (2017), Hateful Memes Kiela et al. (2020), SST2 Radford et al. (2021), ImageNet Deng et al. (2009). Details on each dataset and the corresponding evaluation metrics are provided in Tab. 12.
B.2 Detail Linear Probe Results
Following ALIP, we conduct experiments on randomly selected subsets of 10M and 30M from the LAION400M dataset. For a comprehensive comparison, we report the linear probe performance on 26 downstream datasets. The complete experimental results are shown in Tab.11. RWKV-CLIP-B/32 outperforms ALIP-ViT-B/32 2.6% and 1.4% when training on LAION10M and LAION30M, respectively. Additionally, RWKV-CLIP-B/16 also surpasses ALIP-ViT-B/16 by 2.1% on average across the 26 datasets. These experimental results indicate that RWKV-CLIP demonstrates both robustness and extensibility.
Appendix C More Visualize and Analysis
C.1 Class Activation Map
As shown in Fig. 8, we visualize the class activation maps of ALIP and RWKV-CLIP on different classes from ImageNet. RWKV-CLIP performs superior in aligning the image patches and textual tokens. For example, RWKV-CLIP captures corresponding text semantic entities in images more accurately.
Figure 8: Class activation maps for ALIP and RWKV-CLIP on different classes from ImageNet.
C.2 Cross Modal Alignment Analysis
To evaluate the performance of the cross-modal alignment of RWKV-CLIP, we random select 50 samples from YFCC15M and visualize the cross-modal cosine similarity matrix in Fig. 9. We observe that the diagonal of the RWKV-CLIP matrix is significantly clearer compared to ALIP, indicating that the representations learned by RWKV-CLIP exhibit greater distinctiveness and improved cross-modal alignment capability.
Figure 9: Visualization of modality gaps.
C.3 Case Study
In Fig. 10, we visualize additional generated text using CapsFusion and our proposed framework. The introduction of detection tags enhances semantic information from images, thereby constraining LLMs and significantly reducing hallucinations.
Figure 10: Comparison of generated text using our proposed diverse description generation framework vs. CapsFusion. Hallucinations are highlighted in red, and additional semantic information is highlighted in green.
CIFAR 10 & CIFAR 100
a photo of a {label}. a blurry photo of a {label}. a black and white photo of a {label}. a low contrast photo of a {label}.
a high contrast photo of a {label}. a bad photo of a {label}. a good photo of a {label}. a photo of a small {label}.
a photo of a big {label}. a photo of the {label}. a blurry photo of the {label}. a black and white photo of the {label}.
a low contrast photo of the {label}. a high contrast photo of the {label}. a bad photo of the {label}. a good photo of the {label}.
a photo of the small {label}. a photo of the big {label}.
Food101
a photo of {label}, a type of food.
Caltech101
a photo of a {label}. a painting of a {label}. a plastic {label}. a sculpture of a {label}.
a sketch of a {label}. a tattoo of a {label}. a toy {label}. a rendition of a {label}.
a embroidered {label}. a cartoon {label}. a {label} in a video game. a plushie {label}.
a origami {label}. art of a {label}. graffiti of a {label}. a drawing of a {label}.
a doodle of a {label}. a photo of the {label}. a painting of the {label}. the plastic {label}.
a sculpture of the {label}. a sketch of the {label}. a tattoo of the {label}. the toy {label}.
a rendition of the {label}. the embroidered {label}. the cartoon {label}. the {label} in a video game.
the plushie {label}. the origami {label}. art of the {label}. graffiti of the {label}.
a drawing of the {label}. a doodle of the {label}.
Stanford Cars
a photo of a {label}. a photo of the {label}. a photo of my {label}. i love my {label}!
a photo of my dirty {label}. a photo of my clean {label}. a photo of my new {label}. a photo of my old {label}.
DTD
a photo of a {label} texture. a photo of a {label} pattern. a photo of a {label} thing. a photo of a {label} object.
a photo of the {label} texture. a photo of the {label} pattern. a photo of the {label} thing. a photo of the {label} object.
FGVC Aircraft
a photo of a {label}, a type of aircraft. a photo of the {label}, a type of aircraft.
Flowers102
a photo of a {label}, a type of flower.
Pets
a photo of a {label}, a type of pet.
SUN39
a photo of a {label}. a photo of the {label}.
ImageNet
a bad photo of a {label}. a photo of many {label}. a sculpture of a {label}. a photo of the hard to see {label}.
a low resolution photo of the {label}. a rendering of a {label}. graffiti of a {label}. a bad photo of the {label}.
a cropped photo of the {label}. a tattoo of a {label}. the embroidered {label}. a photo of a hard to see {label}.
a bright photo of a {label}. a photo of a clean {label}. a photo of a dirty {label}. a dark photo of the {label}.
a drawing of a {label}. a photo of my {label}. the plastic {label}. a photo of the cool {label}.
a close-up photo of a {label}. a black and white photo of the {label}. a painting of the {label}. a painting of a {label}.
a pixelated photo of the {label}. a sculpture of the {label}. a bright photo of the {label}. a cropped photo of a {label}.
a plastic {label}. a photo of the dirty {label}. a jpeg corrupted photo of a {label}. a blurry photo of the {label}.
a photo of the {label}. a good photo of the {label}. a rendering of the {label}. a {label} in a video game.
a photo of one {label}. a doodle of a {label}. a close-up photo of the {label}. a photo of a {label}.
the origami {label}. the {label} in a video game. a sketch of a {label}. a doodle of the {label}.
a origami {label}. a low resolution photo of a {label}. the toy {label}. a rendition of the {label}.
a photo of the clean {label}. a photo of a large {label}. a rendition of a {label}. a photo of a nice {label}.
a photo of a weird {label}. a blurry photo of a {label}. a cartoon {label}. art of a {label}.
a sketch of the {label}. a embroidered {label}. a pixelated photo of a {label}. itap of the {label}.
a jpeg corrupted photo of the {label}. a good photo of a {label}. a plushie {label}. a photo of the nice {label}.
a photo of the small {label}. a photo of the weird {label}. the cartoon {label}. art of the {label}.
a drawing of the {label}. a photo of the large {label}. a black and white photo of a {label}. the plushie {label}.
a dark photo of a {label}. itap of a {label}. graffiti of the {label}. a toy {label}.
itap of my {label}. a photo of a cool {label}. a photo of a small {label}. a tattoo of the {label}.
Table 13: Full list of prompts to evaluate the performance of zero-shot classification on 11 visual recognition datasets.